ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

19. History of Mathematics

1991 MS Classification: 01A40, 85-03, 85A04 Lichtenberg, Heiner, Federal Ministry of Finance, Bonn and Berlin, Germany The decoding of the Callippian Cycle in the Gregorian Calendar for the period between 1900 and 2099 The Gregorian Calendar is an adjustable cyclical, lunisolar time-reckoning-system, based on the so-called Callipian Cycle: 76 tropical years = 940 synodical months = 27,759 days (Callippos from Kyzikos, about 330 B. C.). I present the Callipian Cycle as it is valid now. My aim is to establish a formula between the (well-known) solar dates and the (unknown) lunar dates of the Gregorian Calendar. (For a special case, namely the solar date for the 14th day of the lunar month Nisanu in the year X, see (1).) Unfortunately, I have to point out a serious danger for this ingenious system by an unhappy proposal of the World Council of Churches of March 1997.

(1) H. Lichtenberg, Hist. Math. 24(1997), 441. Bonn, April/29/1998 = Aiaru/3/1998.